OpenStudy (anonymous):

Solve the system of equations below and enter your answer as an ordered pair. 6x - 4y = -8 11x + 4y = 76

OpenStudy (anonymous):

Add the two equations. 17x = 68 x = 4 6(4) - 4y = -8 24 - 4y = -8 32 = 4y y = 8 (4, 8)